Following are the demand and supply function of a market

QD A = 8000 - 1000Px, QS A = - 4000+2000Px

i) Find out the market clearing price and quantity.

ii) Plot, on one set of axes, the market demand curve and the market supply curve

for commodity A and show the equilibrium point.

iii) Is the equilibrium stable? Explain.


1
Expert's answer
2021-11-14T17:37:14-0500

(i)

"Q_D=8000-1000P_x"

"Q_S=-4000+2000P_x"

Market clearing price and quantity are found at the point where "Q_D=Q_S"

"8000-1000P_x=-4000+2000P_x"

"\implies P_x=4"

Quantity:

"=8000-1000(4)=4000"

Price: "P_x=4"

(ii)




(iii)

The equilibrium is stable.

This is because the system will gravitate back to initial equilibrium position after a shock is experienced.

The demand demand is negatively sloped while the supply curve is positively sloped.